The Smith family has a pool 25 feet long and 20 feet wide. There is a walkway 4.5 ft wide around the pool. What is the perimeter of the pool with the walkway included.
Answer by JulietG(1812)

Perimeter = 2L + 2W
Perimeter of pool = 25 + 25 + 20 + 20 = 90 feet
Perimeter of the pool + walkway = 29.5 + 29.5 + 24.5 + 24.5 = 108 feet